Burt, a football quarterback, has a pass completion percentage of 55.2%. If the probability that Burt will need 8 or more pass attempts to complete his first pass in a game is at least 0.01, then Burt will be benched and replaced by a backup. Will Burt be benched? Use Excel to find the probability, rounding to three decimal places.
Answer #1

probability that Burt will need 8 or more pass attempts to complete his first pass =P(not make a pass in first 7 attempts)

=(1-0.552)7 =0.004

as this is less than 0.01 therefore Burt will not be benched